OpenStudy (anonymous):

Solve and check for extraneous solutions: \[x=sqrt{28+9x}-4\]

OpenStudy (anonymous):

\[x=\sqrt{28+9x}-4\]

OpenStudy (anonymous):

\[x+4=\sqrt{28+9x} \] \[(x+4)^2=28+9x\] \[x^2+8x+16=28+9x\] \[x^2+8x-9x+16-28=0\] \[x^2-x-12=0\] \(x+3)(x-4)=0\] Either x+3=0 or x-4=0 x=-3 or x=4 Therefore both -3 and 4 are solutions of the given equation.